Solve: (1/5)^(t-2)=125.
Round to 4 decimal places if necessary.



Answer by stanbon(75887)

(1/5)^(t-2)=125
Take the natural log of both sides to get:
(t-2)ln(1/5) = ln 125
t-2= ln125/ln(1/5)
t-2= -3
t=-1
Cheers,
Stan H.